# Generalized Uncertainty Principle from the Regularized Self-Energy

### v2

@@ I Introduction
Phenomenological and experimental implications of the GUP have been investigated in low and high-energy regimes. These include atomic systems Ali et al. (2011); Das and Vagenas (2008), quantum optical systems Pikovski et al. (2012), gravitational bar detectors Marin et al. (2013), gravitational decoherence Petruzziello and Illuminati (2021), gravitational waves Das et al. (2021); Moussa et al. (2021), gravitational tests Scardigli and Casadio (2015), composite particles Kumar and Plenio (2020), astrophysical systems Moradpour et al. (2019), condensed matter systems Iorio et al. (2018), cold atoms Gao and Zhan (2016), macroscopic harmonic oscillators Bawaj et al. (2015), gauge theories Kober (2010), quantum noise Girdhar and Doherty (2020), extended GUP from deformed algebra Segreto and Montani (2022), cosmological models Ashoorioon et al. (2005); Easther et al. (2003); Dabrowski and Wagner (2019); Ali and Majumder (2014); Easther et al. (2001); Ali et al. (2015) and Holography Scardigli and Casadio (2003). Reviews of the GUP, its phenomenology, and its experimental implications can be found in Addazi et al. (2021); Hossenfelder (2013). Our work establishes a connection between T-duality in string theory and the generalized uncertainty principle. Among other things, in this work, we would like to understand if the form of GUP depends on the nature of particles, for example, the spin. It is possible to interpret that the experimental and phenomenological research on GUP provides support in favor of T-duality and string theory.
@@ II Modified Schrödinger–Newton equation
$G(k)\sim-l_{0}^{1/2}\,{\left(k^{2}\right)}^{-3/4}\,e^{-l_{0}\sqrt{k^{2}}}.$ (2)
@@
$=-M\,\int\!\frac{d^{3}k}{{\left(2\pi\right)}^{3}}\;{G_{F}(k)|}_{k^{0}=0}\;\exp\!\left(i\vec{k}\vec{r}\right)$ (4)
$=-\frac{M}{\sqrt{r^{2}+l_{0}^{2}}},$ (5)
@@
$H_{\text{int}}$ $=$ $\int d^{3}r\,V\,T^{00}$ (13)
$=$ $-MG\int_{0}^{2\pi}\int_{0}^{\pi}\int_{0}^{\infty}d^{3}r\frac{\rho(\vec{r})}{\sqrt{r^{2}+l_{0}^{2}}}\,,$ (14)
@@
$d^{3}r=r^{2}\sin^{2}\theta\,d\theta\,d\phi\,dr,$ (16)
@@
$i\hbar\frac{\partial}{\partial t}\Psi(t,\vec{r})=-\frac{\hbar^{2}}{2M}\nabla^{2}\Psi(t,\vec{r})+\hat{H}_{int}\Psi(t,\vec{r}),$ (18)
@@ II.1 Significance of effects in the SN equation
$\Psi(r,t)=\left(\pi\,a^{2}\right)^{-3/4}\,\left(1+\frac{i\,\hbar\,t}{M\,a^{2}}\right)^{-3/2}\,e^{-\frac{r^{2}}{2a^{2}\left(1+\frac{i\,\hbar\,t}{M\,a^{2}}\right)}},$ (23)
@@
$\ddot{r}_{p}(0)=-\frac{d}{dr}\,V(r_{p}(0))=\frac{M\,a}{(l_{0}^{2}+a^{2})^{3/2}}.$ (26)
@@ II.2 Curved spacetime
$ds^{2}=g_{00}\,dt^{2}+g_{rr}\,dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),$ (28)
@@
$\begin{split}f(r)=\left(1-\frac{2Gm(r)}{r}\right)=\left(1-2\Phi(r)\right).\end{split}$ (29)
Using the energy density (15) it was found that
@@ III Regularised self-energy due to the electric field
$H^{em}_{\text{int}}$ $=$ $-Q\int_{0}^{2\pi}\int_{0}^{\pi}\int_{0}^{\infty}d^{3}r\frac{\rho_{em}(\vec{r})}{4\pi\epsilon_{0}\sqrt{r^{2}+l_{0}^{2}}}\,,$ (37)
@@ III.1 Curved spacetime
$f(r)=1-\frac{2Qr^{2}}{{\left(r^{2}+l_{0}^{2}\right)}^{3/2}}=\left(1-2\Phi_{em}(r)\right),$ (39)
@@ IV Modified Klein-Gordon equation
where $\alpha=(9/2)\,\pi^{2}/16^{2}$ and $M_{Pl}=\sqrt{\frac{\hbar c}{G}}$ is the Planck mass. This expression matches the one derived in extended uncertainty principle derived in Mureika (2019). This could be related to gravitational self-completeness with quantum mechanical mass limits. In particular, we may have two cases:
@@ V Modified Dirac equation
$i\gamma^{\mu}\nabla_{\mu}\psi(x,t)-\frac{Mc}{\hbar}\psi(x,t)+\frac{H_{int}}{\hbar c}\psi(x,t)=0,$ (58)
@@
$\psi(x^{\mu})=A\exp\left(-\frac{i}{\hbar}x^{\mu}p_{\mu}\right)\left(\begin{matrix}u_{A}\\ u_{B}\\ \end{matrix}\right)$ (63)
@@ VI Modified Proca equation
$\psi_{\nu}(x^{\mu})=C_{\nu}\exp\left(-\frac{i}{\hbar}x^{\mu}p_{\mu}\right)$ (73)
@@
In other words, the corrections to the energy and mass of the particle with spin 1 particles described by the Proca equation are similar to to scalar particles with spin 0 described by the Klein-Gordon equation.
@@ VII GUP from regularized self-energy
$\left[x_{i},p_{j}\right]=i~{}\delta_{ij}~{}\hbar,$ (84)
@@
$\left[x_{i},p_{j}\right]$ $=$ $i\hbar\left[\delta_{ij}+\beta\delta_{ij}p^{2}+2\beta p_{i}p_{j}\right]$ (85)
@@
$[x_{i},p_{j}]=i\hbar\left[\delta_{ij}-\alpha\left(p\delta_{ij}+\frac{p_{i}p_{j}}{p}\right)\right],~{}~{}~{}$ (86)
@@ VII.1 Quadratic GUP and bosons particles
Let us now show a very interesting result where one can obtain the GUP principle using the modified energy. Let us assume $p<<M_{tot}c^{2}$ in the expressions for corrected energy, we can then rewrite Eqs. (56) and (83) as follows
@@
where $E_{tot}=M_{tot}c^{2}$ and $E=Mc^{2}$. By differentiating with respect to the energy we get
$dE_{tot}=dE\left(1+3\alpha\frac{E^{2}}{c^{4}M_{Pl}^{2}}\right),$ (88)
assuming
$dE_{tot}dt\sim\hbar,$ (89)
@@
$dEdt\sim\hbar\left(1-3\alpha\frac{E^{2}}{c^{4}M_{Pl}^{2}}\right),$ (90)
but since
$l_{0}\sim l_{Pl}=\sqrt{\frac{\hbar G}{c^{3}}}$ (91)
we can write further as
$dEdt\sim\hbar\left(1-\frac{\beta\,l^{2}_{Pl}E^{2}}{c^{2}\hbar^{2}}\right),$ (92)
where $\beta=3\alpha$. This shows that the GUP for the time-energy can be viewed as a consequence of the modified energy of the particle due to the regularized self-energy. To obtain the momentum-time GUP relation, we simply need to use Eqs. (56) and (83) and making use of $\Delta p_{tot}\to cM_{tot}$, along with $\Delta p\to cM$ (similar relations have been used for example in Carr et al. (2015)), and we multiply both sides by $\Delta x$, we get
Finally, if we assume that $\Delta p_{tot}\,\Delta x\sim\hbar$, we obtain
This GUP expression is consistent with the quadratic form of GUP given by Eq. (85), provided $\alpha=-\beta_{0}/2$. In general $\beta_{0}$ is a free parameter which can be positive or negative.
VII.2 Linear GUP and Dirac particles
Consider the corrections to the mass given by Eq. (69) and assume again in the correction for the energy $p<<M_{tot}c^{2}$; we get
$E_{tot}\simeq E\left(1+\alpha\frac{E}{c^{2}M_{Pl}}\right).$ (95)
By differentiating with respect to the energy we get
$dE_{tot}=dE\left(1+2\alpha\frac{E}{c^{2}M_{Pl}}\right),$ (96)
assuming again that
$dE_{tot}dt\sim\hbar,$ (97)
we get
$dEdt\sim\hbar\left(1-2\alpha\frac{E}{c^{2}M_{Pl}}\right),$ (98)
@@ VII.2 Linear GUP and Dirac particles
$dEdt\sim\hbar\left(1-\frac{\beta\,l_{Pl}E}{c\hbar}\right),$ (99)
where $\beta=2\alpha$. This shows again the time-energy GUP may be a consequence of the modified energy of the particle due to the regularized self-energy. If we further take $\Delta p_{tot}\to cM_{tot}$ and $\Delta p\to cM$ in Eq. (69), we obtain
@@
This is the linear GUP in agreement with Eq. (86). We should point out that by comparing to Eq. (86), the correction can be positive or negative. This has to do with the fact that the free parameter $\alpha_{0}$ can be positive or negative. We, therefore, find that the form of GUP depends on the nature of the particles being considered. In particular, for scalar particles with spin zero, we obtained a quadratic form of GUP, while for fermion particles with spin 1/2, we got a linear GUP. This is consistent with the connection between spin and linear GUP that was found in Ali and Majumder (2021) that have implications with explaining the quantum entanglement Ali (2022) through finding a relation between uncertainty and Bekenstein bound Buoninfante et al. (2022).
@@ VIII Conclusions
For the energy we may take $E_{tot}=hf$, with $f$ being the graviton’s frequency, then using the gravitational waves it might be interesting to see whether one can constrain the graviton mass along with the zero-point length. There is also this possibility that we cannot constrain the zero-point length from observations. This has to do with the fact that $M_{tot}^{g}$ might be the true mass being measured by observations and, it simply shifts by some constant compared to the bare mass $M^{g}$, meaning that we cannot distinguish these quantities with observations. We should like to study more about the implication of GUP on the graviton mass in the near future.

### v3

@@ I Introduction
Phenomenological and experimental implications of the GUP have been investigated in low and high-energy regimes. These include atomic systems Ali et al. (2011); Das and Vagenas (2008), quantum optical systems Pikovski et al. (2012), gravitational bar detectors Marin et al. (2013), gravitational decoherence Petruzziello and Illuminati (2021), gravitational waves Feng et al. (2017); Das et al. (2021); Moussa et al. (2021), gravitational tests Scardigli and Casadio (2015); Hemeda et al. (2024), composite particles Kumar and Plenio (2020), astrophysical systems Moradpour et al. (2019), condensed matter systems Iorio et al. (2018), cold atoms Gao and Zhan (2016), macroscopic harmonic oscillators Bawaj et al. (2015), gauge theories Kober (2010), quantum noise Girdhar and Doherty (2020), extended GUP from deformed algebra Segreto and Montani (2022), modified gravity theories Ali and Wojnar (2024), cosmological models Ashoorioon et al. (2005); Easther et al. (2003); Dabrowski and Wagner (2019); Ali and Majumder (2014); Easther et al. (2001); Ali et al. (2015) and Holography Scardigli and Casadio (2003). Reviews of the GUP, its phenomenology, and its experimental implications can be found in Addazi et al. (2021); Hossenfelder (2013). Our work establishes a connection between T-duality in string theory and the generalized uncertainty principle. Among other things, in this work, we would like to understand if the form of GUP depends on the nature of particles, for example, the spin. It is possible to interpret that the experimental and phenomenological research on GUP provides support in favor of T-duality and string theory.
@@ II Modified Schrödinger–Newton equation
$G(k)\sim-l_{0}^{1/2}\,{\left(k^{2}\right)}^{-3/4}\,\mathrm{e}^{-l_{0}\sqrt{k^{2}}}.$ (2)
@@
$=-M\,G\,\int\!\frac{d^{3}k}{{\left(2\pi\right)}^{3}}\;{G_{F}(k)|}_{k^{0}=0}\;\exp\!\left(i\vec{k}\vec{r}\right)$ (4)
$=-\frac{GM}{\sqrt{r^{2}+l_{0}^{2}}},$ (5)
@@
$H_{\text{int}}$ $=$ $\int\mathrm{d}^{3}r\,V\,T^{00}$ (13)
$=$ $-MG\int_{0}^{2\pi}\int_{0}^{\pi}\int_{0}^{\infty}\mathrm{d}^{3}r\frac{\rho(\vec{r})}{\sqrt{r^{2}+l_{0}^{2}}}\,,$ (14)
@@
$\mathrm{d}^{3}r=r^{2}\sin\theta\,\mathrm{d}\theta\,\mathrm{d}\phi\,\mathrm{d}r,$ (16)
@@
$\mathrm{i}\hbar\frac{\partial}{\partial t}\Psi(t,\vec{r})=-\frac{\hbar^{2}}{2M}\nabla^{2}\Psi(t,\vec{r})+\hat{H}_{\text{int}}\Psi(t,\vec{r}),$ (18)
@@ II.1 Significance of effects in the SN equation
$\Psi(r,t)=\left(\pi\,a^{2}\right)^{-3/4}\,\left(1+\frac{\mathrm{i}\,\hbar\,t}{M\,a^{2}}\right)^{-3/2}\,\mathrm{e}^{-\frac{r^{2}}{2a^{2}\left(1+\frac{i\,\hbar\,t}{M\,a^{2}}\right)}},$ (23)
@@
$\ddot{r}_{p}(0)=-\frac{\mathrm{d}}{\mathrm{d}r}\,V(r_{p}(0))=\frac{M\,a}{(l_{0}^{2}+a^{2})^{3/2}}.$ (26)
@@ II.2 Curved spacetime
$\mathrm{d}s^{2}=g_{00}\,\mathrm{d}t^{2}+g_{rr}\,\mathrm{d}r^{2}+r^{2}(\mathrm{d}\theta^{2}+\sin^{2}\theta\,\mathrm{d}\phi^{2}),$ (28)
@@
$\begin{split}f(r)=1-\frac{2Gm(r)}{r}=1+2\Phi(r),\end{split}$ (29)
where $\Phi(r)=-Gm(r)/r$. Using the energy density (15) it was found that
@@ III Regularised self-energy due to the electric field
$H^{em}_{\text{int}}$ $=$ $-Q\int_{0}^{2\pi}\int_{0}^{\pi}\int_{0}^{\infty}\mathrm{d}^{3}r\frac{\rho_{em}(\vec{r})}{4\pi\epsilon_{0}\sqrt{r^{2}+l_{0}^{2}}}\,,$ (37)
@@ III.1 Curved spacetime
$f(r)=1-\frac{2Qr^{2}}{{\left(r^{2}+l_{0}^{2}\right)}^{3/2}}=\left(1+2\Phi_{em}(r)\right),$ (39)
@@ IV Modified Klein-Gordon equation
where $\alpha=(9/2)\,\pi^{2}/16^{2}$ and $M_{Pl}=\sqrt{\frac{\hbar c}{G}}$ is the Planck mass. This expression matches the one derived in the extended uncertainty principle derived in Mureika (2019). This could be related to gravitational self-completeness with quantum mechanical mass limits. In particular, we may have two cases:
@@ V Modified Dirac equation
$\mathrm{i}\gamma^{\mu}\nabla_{\mu}\psi(x,t)-\frac{Mc}{\hbar}\psi(x,t)+\frac{H_{int}}{\hbar c}\psi(x,t)=0,$ (58)
@@
$\psi(x^{\mu})=A\exp\left(-\frac{\mathrm{i}}{\hbar}x^{\mu}p_{\mu}\right)\left(\begin{matrix}u_{A}\\ u_{B}\\ \end{matrix}\right)$ (63)
@@ VI Modified Proca equation
$\psi_{\nu}(x^{\mu})=C_{\nu}\exp\left(-\frac{\mathrm{i}}{\hbar}x^{\mu}p_{\mu}\right)$ (73)
@@
In other words, the corrections to the energy and mass of the particle with spin 1 particles described by the Proca equation are similar to scalar particles with spin 0 described by the Klein-Gordon equation.
@@ VII GUP from regularized self-energy
$\left[x_{i},p_{j}\right]=\mathrm{i}~{}\delta_{ij}~{}\hbar,$ (84)
@@
$\left[x_{i},p_{j}\right]$ $=$ $\mathrm{i}\hbar\left[\delta_{ij}+\beta\delta_{ij}p^{2}+2\beta p_{i}p_{j}\right]$ (85)
@@
$[x_{i},p_{j}]=\mathrm{i}\hbar\left[\delta_{ij}-\alpha\left(p\delta_{ij}+\frac{p_{i}p_{j}}{p}\right)\right],~{}~{}~{}$ (86)
@@ VII.1 Quadratic GUP and bosons particles
Let us show a very interesting result where one can obtain the GUP principle using the modified energy. Let us rewrite Eqs. (56) and (83) as follows
@@
where $E_{tot}=M_{tot}c^{2}$ and $E=Mc^{2}$. By differentiating with respect to the energy on both sides we get
$\mathrm{d}E_{tot}\simeq\mathrm{d}E\left(1+3\alpha\frac{E^{2}}{c^{4}M_{Pl}^{2}}\right).$ (88)
Only in the limit $\alpha\to 0$, for energies we have $E_{tot}=E$, therefore, in general, we expect two possibilities. The first possibility is to assume
$\mathrm{d}E_{tot}\mathrm{d}t\sim\hbar.$ (89)
If we multiply Eq. (87) in both sides with $dt$ we then get in leading order terms
$\mathrm{d}E\mathrm{d}t\sim\hbar\left(1-3\alpha\frac{E^{2}}{c^{4}M_{Pl}^{2}}\right),$ (90)
but since
$l_{0}\sim l_{Pl}=\sqrt{\frac{\hbar G}{c^{3}}},\,\,\,M_{Pl}=\sqrt{\frac{\hbar c}{G}},$ (91)
we can write Eq. (90) further as
$\mathrm{d}E\mathrm{d}t\sim\hbar\left(1-\frac{\beta\,l^{2}_{Pl}E^{2}}{c^{2}\hbar^{2}}\right),$ (92)
where $\beta=3\alpha$. The second possibility is to assume
$\mathrm{d}E\mathrm{d}t\sim\hbar,$ (93)
and, again, by multiplying Eq. (87) with $dt$ we get
$\mathrm{d}E_{tot}\mathrm{d}t\sim\hbar\left(1+\frac{\beta\,l^{2}_{Pl}E^{2}}{c^{2}\hbar^{2}}\right).$ (94)
In other words, there are two ways of expressing the GUP corrected time-energy relation with the difference in the sign. The difference in the sign before the second term is of course related to whether we choose to work with the total or the bare energy.
This shows that the GUP for the time-energy can be viewed as a consequence of the modified energy of the particle due to the regularized self-energy. To obtain the momentum-time GUP relation, we simply need to use Eqs. (56) and (83) and making use of $\Delta p_{tot}\to cM_{tot}$, along with $\Delta p\to cM$ (similar relations have been used for example in Carr et al. (2015)), and we multiply both sides by $\Delta x$, we get
We have a similar situation where $\Delta p_{tot}=\Delta p$, provided $\alpha\to 0$. To this end, we can first assume $\Delta p_{tot}\,\Delta x\sim\hbar$. In that case we obtain
This GUP expression is consistent with the quadratic form of GUP given by Eq. (85), provided $\alpha=-\beta_{0}/2$. The second case is to assume $\Delta p\,\Delta x\sim\hbar$, then from Eq. (95) we obtain
$\Delta p_{tot}\,\Delta x\sim\hbar\,\left(1+\frac{\alpha\,l^{2}_{Pl}\,\Delta p^{2}}{\hbar^{2}}\right).$ (97)
We, therefore, found that in general there are two equivalent representations of the GUP corrected momentum-position relation. In general $\beta_{0}$ is a free parameter that can be positive or negative. This sign in the second term is also related to the fact whether we work with $\Delta p_{tot}$ or $\Delta p$.
VII.2 Linear GUP and Dirac particles
Consider the corrections to the mass given by Eq. (69) which can be written as
$E_{tot}\simeq E\left(1+\alpha\frac{E}{c^{2}M_{Pl}}\right).$ (98)
By differentiating with respect to the energy on both sides we get
$\mathrm{d}E_{tot}=\mathrm{d}E\left(1+2\alpha\frac{E}{c^{2}M_{Pl}}\right).$ (99)
As in the above discussion, let us consider first the case
$\mathrm{d}E_{tot}\mathrm{d}t\sim\hbar,$ (100)
@@ VII.2 Linear GUP and Dirac particles
$\mathrm{d}E\mathrm{d}t\sim\hbar\left(1-2\alpha\frac{E}{c^{2}M_{Pl}}\right),$ (101)
@@
$\mathrm{d}E\mathrm{d}t\sim\hbar\left(1-\frac{\beta\,l_{Pl}E}{c\hbar}\right),$ (102)
where $\beta=2\alpha$. To get the second representation we can assume
$\mathrm{d}E\mathrm{d}t\sim\hbar,$ (103)
in Eq. (98). If follows then
$\mathrm{d}E_{tot}\mathrm{d}t\sim\hbar\left(1+\frac{\beta\,l_{Pl}E}{c\hbar}\right).$ (104)
This shows again that in both cases the time-energy GUP relation can be a consequence of the modified energy of the particle due to the regularized self-energy. If we further take $\Delta p_{tot}\to cM_{tot}$ and $\Delta p\to cM$ in Eq. (69), we obtain
@@
Finally, if we assume $\Delta p\,\Delta x\sim\hbar$ in Eq. (105), we the modified GUP relation
$\Delta p_{tot}\,\Delta x\sim\hbar\,\left(1+\frac{\alpha\,l_{Pl}\,\Delta p}{\hbar}\right).$ (107)
This is the linear GUP in agreement with Eq. (86). We should point out that by comparing it to Eq. (86), the correction can be positive or negative. This has to do with the fact that the free parameter $\alpha_{0}$ can be positive or negative. We, therefore, find that the form of GUP depends on the nature of the particles being considered. In particular, for scalar particles with spin zero, we obtained a quadratic form of GUP, while for fermion particles with spin 1/2, we got a linear GUP. This is consistent with the connection between spin and linear GUP that was found in Ali and Majumder (2021) that have implications with explaining the quantum entanglement Ali (2022) through finding a relation between uncertainty and Bekenstein bound Buoninfante et al. (2022).
@@ VIII Conclusions
For the energy we may take $E_{tot}=hf$, with $f$ being the graviton’s frequency, then using the gravitational waves it might be interesting to see whether one can constrain the graviton mass along with the zero-point length. There is also this possibility that we cannot constrain the zero-point length from observations. This has to do with the fact that $M_{tot}^{g}$ might be the true mass being measured by observations and, it simply shifts by some constant compared to the bare mass $M^{g}$, meaning that we cannot distinguish these quantities with observations. We would like to study more about the implication of GUP on the graviton mass in the near future.